I am no good at working out word problems culd someone please help me with this? I am not sure how to turn it into to an equation r how to solve it. Please help me.
a plane took 1 hour longer to travel 560 miles on the firet portion of a flight then it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?

Answer by checkley71(8403)

DISTANCE=SPEED*TIME
DISTANCE/SPEED=TIME
560/X=(480/X)+1 NOW CROSS MULTIPLY
X[(480/X)+1]=560
480+X=560
X=560-480
X=80 SPEED OF THE PLANE
560/80=7
(480/80)+1=6+1=7